OpenStudy (anonymous):

a 10-ml pipette delivers 9.99, 10.01, 9.97, and 10.08 mL in an experiment. calculate the mean value and the average deviation from the mean

OpenStudy (aaronq):

mean, is another word of average \(\sf mean=\dfrac{9.99+10.01+9.97+ 10.08}{4}=10.0125=10.01~mL\) I'm not entirely sure what they mean by average deviation, but i assume it's \(\sf \Delta deviation=\dfrac{\sum|x_i-mean|}{N}\) \(\sf =\dfrac{|9.99-10.01|+|10.01-10.01|+|9.97-10.01|+|10.08-10.01|}{4}=\dfrac{0.13}{4}\) =0.0325=0.04 so the value is \(\sf10.01\pm0.04~mL\)